Complete the square to make the binomial a perfect square trinomial.
x^2 - 12x

Found 2 solutions by texttutoring, CharlesG2:
Answer by texttutoring(324)

x^2 - 12x

You have to halve the "b" term, -12, then square it: (-12/2)^2 = (-6)^2 = 36

x^2 - 12x + 36 - 36
(x-6)^2 -36

Answer by CharlesG2(834)

Complete the square to make the binomial a perfect square trinomial.
x^2 - 12x
x^2 - 12x = 0 (set it to zero)
x^2 - 12x + (-6)^2 = (-6)^2 (added the square of half of -12 to both sides)
x^2 - 12x + 36 = 36
if you want to solve for x:
(x - 6)^2 = 6^2
x - 6 = +- 6 (this means that -6 * -6 = 36 or 6 * 6 = 36)
x - 6 = 6 OR x - 6 = -6
x = 12 OR x = 0